OpenStudy (anonymous):

Find the inverse of the function y = 0.4x + 4.

OpenStudy (aravindg):

Try to write x in terms of y.

OpenStudy (anonymous):

Would x= 4.4-y be right?

OpenStudy (aravindg):

Try again?

OpenStudy (anonymous):

Okay, so how do I do this. I have no idea, honestly.

OpenStudy (aravindg):

y = 0.4x + 4. subtract 4 on both sides

OpenStudy (anonymous):

okay, so I get y-4=0.4x ?

OpenStudy (aravindg):

yep

OpenStudy (aravindg):

Now divide by 0.4 on both sides.

OpenStudy (anonymous):

so... (y-4)/0.4 = x?

OpenStudy (aravindg):

yes now replace x with f-1 .You are done!

OpenStudy (anonymous):

Woohoo! Thank you so much!!!

OpenStudy (aravindg):

yw :)